OpenStudy (anonymous):

In a circuit, suppose L = 5, R = 25, E(t) = 100cos60t, E(0) = 100, I(0) = 0. Find I(t).

OpenStudy (espex):

\[I(t) = i_0 * e^{-(\frac{R}{L})t}, t \ge 0\]

OpenStudy (anonymous):

I understand that I(t) = Isp(t) + Itr(t). I'm just struggling with finding these two equations.

OpenStudy (espex):

Are you doing this as an ODE or actually solving the circuit?

OpenStudy (anonymous):

I'm interested in solving it using an ODE.

OpenStudy (espex):

What tools have you developed thus far for solving ODEs?

OpenStudy (espex):

Looking back through my notes it seems that we assumed a series circuit and used an integrating factor to solve. So I may not have the expertise you need. Perhaps @Callisto or @saifoo.khan can help.